OpenStudy (anonymous):

[10.04] What is the exact value of the expression 5the square root of 13. + the square root of 6. - 7the square root of 13.? -2the square root of 13. + the square root of 6. -2the square root of 6. -2the square root of 13. + 2the square root of 6. -the square root of 19.

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 5\sqrt{13}+\sqrt{6}&-7\sqrt{13}\\ \quad \uparrow &\quad \uparrow \\ \end{array}\) add like terms

OpenStudy (anonymous):

then what

OpenStudy (jdoe0001):

then append the \(\bf \sqrt{6}\) to that

OpenStudy (anonymous):

you add them

OpenStudy (jdoe0001):

yeap, when the radicand is the same, you just treat them like same-variables, and add them

OpenStudy (anonymous):

ok then what do you do next

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 5\sqrt{13}+\sqrt{6}&-7\sqrt{13}\\ \quad \uparrow &\quad \uparrow \\ \end{array}\implies \begin{array}{cllll} (5\sqrt{13}-7\sqrt{13})+\sqrt{6}\\ \uparrow \\ \square \sqrt{13}+\sqrt{6} \end{array}\)

OpenStudy (jdoe0001):

nothing else really, that'd simplify it

OpenStudy (anonymous):

A or C what is the answer

OpenStudy (jdoe0001):

well, add them what do you get?

OpenStudy (anonymous):

add what

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 5\sqrt{13}+\sqrt{6}&-7\sqrt{13}\\ \quad \uparrow &\quad \uparrow \\ \end{array}\) the like terms

OpenStudy (anonymous):

-2